# On the Robustness of RSMA to Adversarial BD-RIS-Induced Interference

### v2

@@ IV-C Complexity Analysis
The analysis for the random strategy shows that BD-RIS impose a greater computational load compared to a conventional single-connected RIS, which has a linear complexity of $\mathcal{O}(D)$. For the aligned attack with single-connected RIS, iterative optimization methods are often employed [6, 8], which typically have lower per-iteration complexity but may require multiple iterations for convergence. In contrast, Algorithms 1 and 2 are non-iterative, ensuring a feasible runtime. The added cost of BD-RIS represents a trade-off for their superior adversarial performance, as will be seen in Section V.
@@ V Numerical Results
The simulation setup comprises: (i) a BS equipped with $M=32$ transmit antennas positioned at the origin; (ii) three users located at distances of $d_{1}=30$ m, $d_{2}=50$ m, and $d_{3}=60$ m from the BS, with azimuth angles of $\theta_{1}=25^{\circ}$, $\theta_{2}=15^{\circ}$, and $\theta_{3}=10^{\circ}$; and (iii) an adversarial RIS deployed at a distance of $d_{\text{RIS}}=40$ m from the BS with an azimuth angle of $\theta_{\text{RIS}}=5^{\circ}$. Unless stated otherwise, simulations are performed for $D=200$ reflecting elements, with the group-connected RIS configured as $D_{1}=\cdots=D_{G}=5$, considering a standard imperfect CSI scenario with $\varepsilon^{\text{\tiny BS-U}}=\varepsilon^{\text{\tiny BS-RIS}}=\varepsilon^{\text{\tiny RIS-U}}=0.3$. The small-scale fading for all channel links is modeled following a Rayleigh distribution, representing a rich scattering environment, combined with a power-law path loss model $d^{-\eta}$ with $\eta=3$, where $d$ represents the corresponding distance. In particular, the distance from the RIS to the $u$-th user is obtained by $d_{\text{RIS},u}=\sqrt{d_{\text{RIS}}^{2}+d_{u}^{2}-2d_{\text{RIS}}d_{u}\cos(\theta_{\text{RIS}}-\theta_{u})}$, resulting in $d_{\text{RIS},1}\approx 15.64$ m, $d_{\text{RIS},2}\approx 12.68$ m, and $d_{\text{RIS},3}\approx 20.45$ m. Moreover, the user weights for Algorithms 1 and 2 are set uniformly as $\mu_{1}=\mu_{2}=\mu_{3}=1/3$, the noise power is set to 60-60 dBm, SDMA employs a uniform power allocation with regularized zero-forcing precoders, and RSMA operates under perfect SIC by default. The sensitivity to SIC imperfections, as well as the impact of varying CSI errors, is investigated in Section V-D.
@@ V-A Impact of the RIS Uplink Mode
Fig. 3 investigates the impact of the RIS operation mode during the uplink training phase on the system average sum rate under aligned interference attacks. Fig. 3(a) evaluates this impact across different RIS architectures for the RSMA scheme. It can be seen that using a reflective mode during uplink training significantly boosts the degradation potential for all architectures, but the benefit for the attacker is largest when using BD-RIS. The fully connected RIS, for instance, reduces the sum rate to $23.4$ bits/s/Hz at $40$ dBm, a more severe degradation compared to the $24.7$ bits/s/Hz achieved by the single-connected RIS.
Fig. 3(b) then provides a direct comparison between RSMA and SDMA for the most potent fully connected architecture. The results confirm that the reflective mode is more damaging than the absorption mode for both multiple access schemes. The figure also sheds light on RSMA’s inherent robustness, as it consistently outperforms SDMA under all attacked scenarios. This superior performance is a result of the adaptive power allocation described in Section III-F, which allows RSMA to better manage the channel estimation errors induced by the reflective mode attack. These observations provide the first evidence of RSMA’s robustness advantage over SDMA, which will be further analyzed in the next subsection using the proposed robustness metrics.
@@ V-B Robustness of RSMA
Complementing the insights above, Fig. 4(b) quantifies the sum-rate degradation of RSMA across different RIS architectures. For all attack types, the degradation curves peak near $12$ dBm and then decrease slightly as transmit power increases, a behavior resulting from RSMA’s flexible interference management mechanism. The figure also confirms that the aligned attack with a fully connected RIS yields the largest degradation among the considered architectures.
V-C Impact of the Number of RIS Reflecting Elements
Fig. 5 analyzes the scaling behavior of the adversarial RIS for a varying number of reflecting elements $D$ under the aligned interference attack, considering a fixed transmit power of $30$ dBm. As shown in Fig. 5(a), the average sum rate for both multiple access schemes degrades as the RIS becomes larger. The performance of SDMA, however, collapses at large $D$, particularly when attacked by a fully connected RIS, whereas the decline of RSMA is considerably less severe. For instance, at $D=400$, SDMA falls below $12$ bits/s/Hz, while RSMA still sustains above $16$ bits/s/Hz when both are attacked by the fully connected RIS.
@@ V-C Impact of the Number of RIS Reflecting Elements
Figure 5: Impact of the number of reflecting elements $D$ for different RIS architectures: (a) average sum rate; and (b) average sum-rate degradation. Results assume a fixed transmit power of $30$ dBm under the aligned interference attack.
@@ V-D Impact of CSI and SIC Imperfections
Fig. 6 provides a deeper analysis of the impact of imperfect CSI, considering a fixed transmit power of $30$ dBm for both random and aligned attacks induced by the fully connected RIS. Fig. 6(a) investigates system performance versus the error in the legitimate BS-user link $\epsilon^{\text{\tiny BS-U}}$, while the attacker’s CSI for the RIS-related links has a fixed error of $\epsilon^{\text{\tiny BS-RIS}}=\epsilon^{\text{\tiny RIS-U}}=0.3$. As can be seen, at $\epsilon^{\text{\tiny RIS-U}}=0$, RSMA and SDMA exhibit identical performance under the same attack. This behavior is explained by the fact that RSMA reduces to SDMA when the BS has perfect channel knowledge. As the error in the legitimate channel increases, the sum rate of SDMA under both attack types monotonically decreases. RSMA, in contrast, shows a slight initial increase in its sum rate for small $\epsilon^{\text{\tiny BS-U}}>0$ under both random and aligned attacks.
@@
Figure 7: Impact of imperfect SIC on the system sum rate: (a) versus transmit power for different SIC error factors $\xi$; and (b) versus the SIC error factor $\xi$ for a fixed transmit power of $30$ dBm. Results assume the aligned interference attack induced by the fully connected RIS.
@@
The impact of the SIC error factor is further detailed in Fig. 7(b) for a fixed transmit power of $30$ dBm. The performance of RSMA shows a steep decline as $\xi$ increases, for both the safe and attacked scenarios. In particular, RSMA approaches the performance of SDMA from $\xi=0.6\times 10^{-3}$ onward. On one hand, this exposes the sensitivity of RSMA to imperfect SIC. On the other hand, it reveals the effectiveness of its adaptive power allocation, which, being aware of SIC errors, reallocates more power to private messages so that RSMA at least matches SDMA under imperfect SIC. This finding highlights a critical trade-off: while RSMA is fundamentally more robust to external CSI uncertainties, its overall performance advantage is contingent on the availability of high-quality hardware for interference cancellation.
@@ VI Conclusions
This paper has investigated the robustness of RSMA-based multi-user MISO systems to adversarial attacks induced by BD-RIS. Two effective attack strategies were described and studied: random and aligned interference. For each, novel algorithms were proposed to generate valid reflection coefficients respecting the architectural BD-RIS constraints. Extensive simulations revealed that RSMA demonstrated significantly greater robustness to these attacks than the benchmark SDMA scheme, especially under imperfect CSI. Moreover, BD-RIS significantly enhances adversarial efficacy compared to conventional RIS, with particularly severe performance degradation observed when the BD-RIS manipulates the uplink training phase via reflective operation. Future work should expand this analysis and investigate the adversarial potential of multi-sector BD-RIS and its non-reciprocal architectures. The development of robust countermeasures against such threats is another crucial direction.
@@ Acknowledgments
This work was supported by the Commonwealth Cyber Initiative (cyberinitiative.org) in Virginia, US, an investment to advance cyber R&D, innovation, and workforce development, the National Science Foundation under grants no. 2326599 and 231879, and the Research Council of Finland through the 6G-ConCoRSe project (grant no. 359850) and the 6G Flagship programme (grant no. 369116).

### v3

@@ IV-C Complexity Analysis
The analysis for the random strategy shows that BD-RIS impose a greater computational load compared to a conventional single-connected RIS, which has a linear complexity of $\mathcal{O}(D)$. For the aligned attack with single-connected RIS, iterative optimization methods are often employed [8, 41], which typically have lower per-iteration complexity but require multiple iterations for convergence. In contrast, Algorithms 1 and 2 are non-iterative, and their computational cost is primarily driven by the dimensionality of the considered BD-RIS architectures. To illustrate this behavior, Table I reports the average runtime of the considered attack strategies for different numbers of RIS elements. The results show that the runtime of all methods increases with the RIS size, in agreement with the complexity scaling trends discussed earlier. The single-connected RIS baseline based on the iterative optimization in [8] presents higher runtime values up to $D=200$. However, its lower per-iteration complexity enables it to become more competitive for larger RIS sizes (i.e., $D\geq 250$), despite its iterative nature. In this large-scale regime, the fully connected RIS becomes the most computationally demanding configuration, consistent with its quartic scaling in $D$. On the other hand, the group-connected RIS (with a group size fixed to five elements) exhibits the lowest runtime values across all evaluated configurations, reflecting the reduced problem dimension achieved by its partitioned architecture.
The above analysis provides insight into the practical implications of the proposed attacks. On the one hand, the worst-case complexity of $\mathcal{O}(UMD^{4})$ for the fully connected RIS, which is reflected in the runtimes reported in Table I, indicates that scaling fully connected architectures to very large sizes may become challenging. On the other hand, the group-connected RIS significantly reduces the computational burden, thus offering a more scalable BD-RIS implementation, while still causing a strong adversarial impact, as will be seen in Section V.
TABLE II: Summary of simulation parameters
Parameter Value
Number of BS antennas $M$ $32$
Number of RIS elements $D$ $200$
Group-connected RIS size $D_{g}$ 55
Number of users $U$ 33
User distances $\{d_{u}\}$ $\{30,50,60\}$ m
User azimuth angles $\{\theta_{u}\}$ $\{25^{\circ},15^{\circ},10^{\circ}\}$
RIS distance $d_{\text{RIS}}$ $40$ m
RIS azimuth angle $\theta_{\text{RIS}}$ $5^{\circ}$
Path-loss exponent $\eta$ 33
CSI error factors $\epsilon^{\text{\tiny BS-U}},\epsilon^{\text{\tiny BS-RIS}},\epsilon^{\text{\tiny RIS-U}}\in[0,0.7]$
SIC error factor $\xi\in[0,7\times 10^{-3}]$
Noise power 60-60 dBm
@@ V Numerical Results
The simulation setup comprises: (i) a BS equipped with $M=32$ transmit antennas positioned at the origin; (ii) three users located at distances of $d_{1}=30$ m, $d_{2}=50$ m, and $d_{3}=60$ m from the BS, with azimuth angles of $\theta_{1}=25^{\circ}$, $\theta_{2}=15^{\circ}$, and $\theta_{3}=10^{\circ}$; and (iii) an adversarial RIS deployed at a distance of $d_{\text{RIS}}=40$ m from the BS with an azimuth angle of $\theta_{\text{RIS}}=5^{\circ}$. Unless stated otherwise, simulations are performed for $D=200$ reflecting elements, with the group-connected RIS configured as $D_{1}=\cdots=D_{G}=5$, considering a standard imperfect CSI scenario with $\epsilon^{\text{\tiny BS-U}}=\epsilon^{\text{\tiny BS-RIS}}=\epsilon^{\text{\tiny RIS-U}}=0.3$. The small-scale fading for all channel links is modeled following a Rayleigh distribution, representing a rich scattering environment, combined with a power-law path loss model $d^{-\eta}$ with $\eta=3$, where $d$ denotes the distance. In particular, the distance from the RIS to the $u$-th user is obtained by $d_{\text{RIS},u}=\sqrt{d_{\text{RIS}}^{2}+d_{u}^{2}-2d_{\text{RIS}}d_{u}\cos(\theta_{\text{RIS}}-\theta_{u})}$, resulting in $d_{\text{RIS},1}\approx 15.64$ m, $d_{\text{RIS},2}\approx 12.68$ m, and $d_{\text{RIS},3}\approx 20.45$ m. Moreover, the user weights for Algorithms 1 and 2 are set uniformly as $\mu_{1}=\mu_{2}=\mu_{3}=1/3$, the noise power is set to 60-60 dBm, SDMA employs a uniform power allocation with regularized zero-forcing precoders, and RSMA operates under perfect SIC by default. Table II summarizes simulated parameter values. The sensitivity to SIC imperfections, as well as the impact of varying CSI errors, is investigated in Section V-D.
@@ V-A Impact of the RIS Uplink Mode
Fig. 4 investigates the impact of the RIS operation mode during the uplink training phase on the system average sum rate under aligned interference attacks. Fig. 4(a) evaluates this impact across different RIS architectures for the RSMA scheme. It can be seen that using a reflective mode during uplink training significantly boosts the degradation potential across all architectures, but the benefit to the attacker is greatest when using BD-RIS. The fully connected RIS, for instance, reduces the sum rate to $23.4$ bits/s/Hz at $40$ dBm, a more severe degradation compared to the $24.7$ bits/s/Hz achieved by the single-connected RIS.
Fig. 4(b) then provides a direct comparison between RSMA and SDMA for the most potent fully connected architecture. The results confirm that the reflective mode is more damaging than the absorption mode for both multiple access schemes. The figure also sheds light on RSMA’s inherent robustness, as it consistently outperforms SDMA across all attack scenarios. This superior performance is a result of the adaptive power allocation described in Section III-F, which enables RSMA to better manage channel estimation errors induced by the reflective mode attack. These observations provide the first evidence of RSMA’s robustness advantage over SDMA, which will be further analyzed in the next subsection using the proposed robustness metrics.
@@
Figure 6: Impact of RIS size and architecture parameters: (a) average sum rate versus $D$; (b) average sum-rate degradation versus $D$; and (c) average sum rate versus transmit power for different group sizes $D_{g}$ of the group-connected RIS. In (a) and (b), the group-connected RIS uses the default group size $D_{g}=5$ and a fixed transmit power of $30$ dBm. Results assume the aligned interference attack.
@@ V-B Robustness of RSMA
Another important characteristic of the robustness curves in Fig. 5(a) is the non-monotonic trend of RSMA, with a noticeable rebound after $12$ dBm. The origin of this behavior becomes clear when jointly examining the robustness curves and the corresponding sum-rate degradations of RSMA across different RIS architectures in Fig. 5(b). The degradation curves directly quantify the relative gap between the sum rates of the attacked and safe systems, i.e., $\sum_{\forall u}\Delta R_{u}$, as defined in (10). In particular, for all attack types, the degradation curves peak near $12$ dBm and then decrease gradually as the transmit power increases. This behavior is directly reflected in the robustness index defined in (11), which normalizes the rate loss under attack by the corresponding safe-system rate. Consequently, the peak degradation around $12$ dBm corresponds to the local minimum observed in the robustness index curves, while the subsequent reduction in degradation at higher transmit powers leads to the observed rebound of RSMA. This rebound is consistent with the adaptive interference-management mechanism of RSMA: as the impact of channel acquisition mismatch becomes more pronounced, the power allocation shifts toward the common message, mitigating the impact of inaccurate CSI and reducing the relative rate gap to the safe case. In contrast, SDMA lacks this adaptive mechanism, resulting in a more steadily decreasing robustness index under the same attack conditions. Finally, Fig. 5(b) also confirms that the aligned attack with a fully connected RIS yields the largest degradation among the considered architectures, which is consistent with the corresponding lower robustness levels observed in Fig. 5(a).
V-C Impact of RIS Architecture Parameters
Figs. 6(a) and 6(b) analyze the scaling behavior of the adversarial RIS for a varying number of reflecting elements $D$ under the aligned interference attack, considering a fixed transmit power of $30$ dBm, while Fig. 6(c) examines the impact of the group size $D_{g}$ in the group-connected RIS for a fixed number of reflecting elements. As shown in Fig. 6(a), the average sum rate for both multiple access schemes degrades as the RIS becomes larger. The performance of SDMA, however, collapses at large $D$, particularly when attacked by a fully connected RIS, whereas the decline of RSMA is considerably less severe. For instance, at $D=400$, SDMA falls below $12$ bits/s/Hz, while RSMA still sustains above $16$ bits/s/Hz when both are attacked by the fully connected RIS.
@@ V-C Impact of RIS Architecture Parameters
Fig. 6(c) shows the effect of the group size $D_{g}$ in the group-connected RIS, highlighting the tradeoff between implementation complexity and attack performance. For a fixed number of $D=200$ reflecting elements, increasing $D_{g}$ leads to a stronger attack and increased sum-rate degradation. Smaller group sizes result in lower performance and behavior closer to that of a single-connected RIS, while benefiting from reduced hardware complexity and lower algorithmic complexity in the attack optimization, as discussed in Section IV-C. In contrast, larger group sizes progressively approach the performance of a fully connected architecture at the cost of increased implementation and optimization complexity. These results confirm that the group-connected RIS provides a flexible architecture that bridges the gap between single-connected and fully connected RIS designs, allowing the adversary to balance complexity and attack effectiveness.
@@ V-D Impact of CSI and SIC Imperfections
Fig. 7 provides a deeper analysis of the impact of imperfect CSI, considering a fixed transmit power of $30$ dBm for both random and aligned attacks induced by the fully connected RIS. Fig. 7(a) investigates system performance versus the error in the legitimate BS-user link $\epsilon^{\text{\tiny BS-U}}$, while the attacker’s CSI for the RIS-related links has a fixed error of $\epsilon^{\text{\tiny BS-RIS}}=\epsilon^{\text{\tiny RIS-U}}=0.3$. As can be seen, at $\epsilon^{\text{\tiny RIS-U}}=0$, RSMA and SDMA exhibit identical performance under the same attack. This behavior is explained by the fact that RSMA reduces to SDMA when the BS has perfect channel knowledge. As the error in the legitimate channel estimates increases, the sum rate of SDMA under both attack types monotonically decreases. RSMA, in contrast, shows a slight initial increase in its sum rate for small $\epsilon^{\text{\tiny BS-U}}>0$ under both random and aligned attacks.
@@
[…]
@@
[…]
Note that SDMA was adopted as the baseline since it represents the dominant multi-user transmission strategy in current MIMO (MIMO) systems and provides a natural benchmark for RSMA, as they share the same spatial multiplexing strategy for private streams. Other multiple access schemes, such as OMA (OMA) and NOMA (NOMA), would also be affected by adversarial RIS-induced imperfect CSI. However, the detailed impact on these schemes depends on their specific transmission and receiver structures. For instance, in works that combine NOMA with MIMO systems, it is common to implement user clustering strategies and multiplex multiple user groups in the spatial domain. In such MIMONOMA systems, adversarial RIS-induced channel manipulation could lead to increased inter-group interference. In OMA-based MIMO systems, where users are orthogonal in time or frequency, they can be served with dedicated precoders, and RIS-induced attacks would impact system performance differently. In this case, the RIS-corrupted CSI would primarily degrade the alignment between the precoder and the true user channels, consequently reducing the BS beamforming gain. A detailed investigation of these effects is beyond the scope of this work and is left for future research.
@@ VI Conclusions
This paper has investigated the robustness of RSMA-based multi-user MISO systems to adversarial attacks induced by BD-RIS. Two effective attack strategies were described and studied: random and aligned interference. For each, novel algorithms were proposed to generate valid reflection coefficients respecting the architectural BD-RIS constraints. Extensive simulations revealed that RSMA demonstrated significantly greater robustness to these attacks than the benchmark SDMA scheme, especially under imperfect CSI. Moreover, BD-RIS significantly enhances adversarial efficacy compared to conventional RIS, with particularly severe performance degradation observed when the BD-RIS manipulates the uplink training phase via reflective operation.
The RSMA system considered in this work is modeled as being unaware of the presence of BD-RIS-adversarial manipulation. Future work should expand this analysis and investigate the development of countermeasures against such threats to further enhance RSMA performance. For example, a robust defense could involve the detection of RIS-induced attacks or anomalies based on uplink-downlink reciprocity checks and statistical monitoring of channel estimates. Once a threat is detected, the BS would proactively adapt the RSMA protocol by adjusting the rate-splitting structure and power allocation, as well as re-optimizing beamforming and channel estimation, thus mitigating the impact of adversarial attacks. In parallel, complementary countermeasures could rely on making the uplink channel estimation phase less predictable through aperiodic or asynchronous training (similar to, e.g., [6]), thus increasing the synchronization burden for the adversary. In addition, further investigation of the adversarial potential of other emerging RIS concepts, such as multi-sector BD-RIS and non-reciprocal architectures, is also an interesting research direction.
@@ Acknowledgments
This work was supported by the Commonwealth Cyber Initiative (cyberinitiative.org) in Virginia, US, an investment to advance cyber R&D, innovation, and workforce development, the National Science Foundation under grants no. 2326599 and 2318798, and the Research Council of Finland through the 6G-ConCoRSe project (grant no. 359850) and the 6G Flagship programme (grant no. 369116).